OpenStudy (anonymous):

Need help with slope? Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4).

OpenStudy (kobeni-chan):

Point slope form is y-y=m(x-x) You fill in each second variable with the coordinate, and m represents slope. x is 3, y is -4, and slope is -3, in this scenario. y+4 = -3(x-3)

OpenStudy (anonymous):

Ok, i was confused because it has been awhile since i have done slope.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ -4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -3 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

well.... no need to solve for "y" in this case... since you need in point-slope form anyhow